Jim is 3 ft from a lamppost that is 15 ft tall. He is 5.5 ft tall. His shadow is ____ ft long. Round your answer to the nearest hundreth.
Answer by Cromlix(4381)

Similar Triangles
Height of Lampost/Jim's height
= Distance between Jim and lampost + x (shadow)/ Distance between
Jim and lampost
15/5.5 = 3 + x/x
Cross multiply
15x = 5.5(3 + x)
15x = 16.5 + 5.5x
15x - 5.5x = 16.5
9.5x = 16.5
x = 1.74ft
His shadow is 1.74ft long
Hope this helps.
:-)